OpenStudy (anonymous):

15i-10i^2 i need help simplifying further please!

OpenStudy (anonymous):

this makes me scared :(

OpenStudy (anonymous):

why?!

OpenStudy (anonymous):

Use this \(i^2=-1\).

OpenStudy (anonymous):

15i-10-1 = 15i-11 am i going in the right direction? :(

OpenStudy (anonymous):

15i-10i^2 =15i-10*(-1) =15i+10

OpenStudy (anonymous):

okay thanks! is this simplest form?

OpenStudy (anonymous):

yes it is.